OpenStudy (anonymous):

The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

zepdrix (zepdrix):

Hey Ariana :)\[\Large\rm d=ut+\frac{1}{2}at^2\]So we need to solve for a, huh?

zepdrix (zepdrix):

We have some addition happening, so we'll need to apply subtraction as our first step. We'll subtract ut from each side,\[\Large\rm d-ut=\frac{1}{2}at^2\]Ok with that step?

OpenStudy (anonymous):

yes i am @zepdrix thanks :D

OpenStudy (javk):

\[d=ut + \frac{ 1 }{ 2 }at ^{2}\] take ut to the other side \[d-ut = \frac{ 1 }{ 2 }at ^{2}\] multiply the entire eqn with 2 to get rid of half \[2d-2ut = at ^{2}\] take t sqrd to the other side \[a=\frac{ 2d-2ut }{{ t}^{2}} \]

OpenStudy (anonymous):

thanks sooo much :D @Javk